A 3.3 kg block slides down a 6.1 m long ramp with a 30° angle of inclination. If the coefficient of kinetic friction between the block and the ramp is 0.49, what is the work done on the block by friction as it slides all the way down?



1
Expert's answer
2021-06-30T11:43:47-0400

Frictional force on block = "\mu" mg cos"\theta"

Here,

"\mu" = 0.49

m = 3.3

g = 10

"\theta" = 30o

Hence, frictional force = "0.49 \times 3.3 \times 10 \times cos(30^o)"

= 14.003 N

And,

mg sin "\theta" = "3.3 \times 10 \times sin(30^o)"

= 16.5 N

Since, frictional force is greater than the driving force, the block will not move.

Hence, the net work done by friction on the block is zero.